Please help me with this problem: The lines 3x + 4y = 5 and 2kx - 5y = 3 are parallel. What is the value of k?
Answer by nerdybill(7384)

Please help me with this problem: The lines 3x + 4y = 5 and 2kx - 5y = 3 are parallel. What is the value of k?
.
If two lines are parallel their slopes are equal.
.
Put both equations into the "slope-intercept" form to discover the slope of each:
3x + 4y = 5
4y = -3x + 5
y = (-3/4)x + 5/4
.
2kx - 5y = 3
-5y = -2kx + 3
y = (2k/5)x - 3/5
.
Set slopes equal to each other:
-3/4 = 2k/5
cross multiply
-15 = 8k
-15/8 = k